OpenStudy (anonymous):

Starting at time t=0, an object moves along a straight line with velocity in m/s given by: v(t) = 98-2t^2 where t is in seconds. When it momentarily stops,what is its acceleration?

OpenStudy (nincompoop):

what is the initial velocity at t=0?

OpenStudy (nincompoop):

solve for t with that formula, which is typically in seconds then obtain the derivative of it your formula then plug the seconds in the derivative to solve for velocity

OpenStudy (anonymous):

@nincompoop you mean like solving the equation first? , well , t = +/-7.

OpenStudy (nincompoop):

ya

OpenStudy (anonymous):

what should I do after this?

OpenStudy (amoodarya):

1. t=7 is correct 2 . put t=7 in v'(t)=d/dt (v(t))=0-4t so a(7)=-4(7) m/s^2

OpenStudy (anonymous):

Thank you @amoodarya